1.
The disposable income of Mehta family increases from Rs 5000 to Rs 15,000. As a result, the family‘s demand for milk and milk goods has increased from 30 liters to 60 liters per month. Calculate the income elasticity of demand.
2.
A drop in the price of lemons from Rs 100 per kg to Rs 60 Per Kg increases the quantity demanded from 1.75 to 7 kg per week. Calculate the price elasticity of demand.

Answer by Theo(13342)

number 1:

the disposable income increases from 5000 to 15000.
the demand for milk and milk goods has increased from 30 liters to 60 liters per month.

percent change in demand = (60 - 30) / 30 = 30 / 30 = 1 * 100 = 1005.

percent change in income = (15000 - 5000) / 5000 = 10000 / 5000 = 2 * 100 = 200%.

the price elasticity of demand = percent change in demand divided by percent change in income = 100 / 200 = .5

what this says is that the percent increase in demand will be 50% of the percent increase in income.

number 2:

percent change in demand = 7 - 1.75 = 5.25 / 1.75 = 3 * 100 = 300%.

percent change in price = 60 - 100 = -40 / 100 = -.4 * 100 = -40%.%

price elasticity of demand = percent change in demand divided by percent change in price = 300 / -40 = -7.5.

what this says is that, if the price goes up 40%, the demand goes down 750%.

here's a reference on elasticity of demand that might be helpful.

https://www.economicshelp.org/blog/7019/economics/examples-of-elasticity/

check my figures to see if they're accurate.
i think they are, but i just learned about what elasticity of demand means and i made my figures based on what the reference was saying.

here's another reference that goes into it a bit further.

https://investinganswers.com/dictionary/e/elasticity